In order to join a voice lesson class you pay a 1,500 pesos fee, then 500 for each class you go to what is the average cost per class graph the function to show whether it form a straight line or a curve
1
Expert's answer
2021-10-28T11:45:26-0400

Let "x=" the number of classes.

The total cost for "x" classes is


"C(x)=1500+500x, x\in \N"

The average cost per class is


"C_{ave}=\begin{cases}
 1500 &x=0 \\
 \dfrac{1500}{x}+500 &x\in\Z^+
\end{cases}"